# Judge Q: Trainable Queries for Optimized Information Retention in KV Cache Eviction

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a wide range of domains, including language modeling (Hendrycks et al. 2020; Suzgun et al. 2022), text comprehension (Rajpurkar et al. 2016; Joshi et al. 2017), and code understanding (Chen et al. 2021; Jain et al. 2024). Currently, many scenarios (Liu et al. 2025a, b; Chen et al. 2025) involve input sequences with over ten thousand tokens, such as books, code repositories (Kočiskỳ et al. 2018; Zhang et al. 2023a), and dialogue systems with long-term histories (Li et al. 2024a). LLMs employ KV cache to store historical input information during the pre-filling stage, and reuse them during the decoding phase. The size of KV cache increases linearly with the length of input sequence. In long sequence scenarios, the KV cache often consumes substantial memory space, posing a significant bottleneck for inference, particularly in resource-limited environments.
@@
This straightforward approach of directly selecting the last window is simple to implement, however, it has limitations. Firstly, it tends to overly focus on local information within the last window while neglecting global content, which can result in potential information loss. Additionally, if the last window does not contain the question, the performance of these methods often significantly degrades. In fact, the primary objective of KV cache eviction is to reduce memory usage while maintaining decoding quality. Therefore, identifying the key-value pairs that are crucial for decoding the subsequent tokens is a logical and ideal KV cache eviction approach.
In this paper, we discover that directly using the original decoded tokens to select critical key-value pairs can maintain a high level of decoding quality, which represents the theoretical upper bound of KV cache eviction methods. However, during the prefilling stage, the actual decoded tokens are not available. Inspired by this phenomenon, we propose a training method called Judge Q. This approach involves introducing a list of soft tokens into the model’s vocabulary and leveraging them to perform KV cache eviction during the pre-filling stage. Specifically, during the training phase, we concatenate these soft tokens directly to the end of original input sequence. The soft tokens are trained to align the attention map to input sequence with that of the actual decoded tokens, enabling the soft tokens to capture global information effectively. This process incurs minimal training costs, as it involves training only the parameters associated with the soft tokens in the model’s embedding layer, while keeping the rest of the model’s weights frozen. During inference, we similarly append these soft tokens to the end of the original input sequence in the pre-filling stage. The queries corresponding to soft tokens are used to calculate the importance scores of the key-value pairs in KV cache across the entire input sequence. Based on these scores, we retain the top-k key-value pairs with the highest scores and discard the rest. After the pre-filling stage, the soft tokens are also removed. The pruned KV cache is then used for the subsequent decoding process. We conduct experiments using two models, Llama-3.1-8B-Instruct (Dubey et al. 2024) and Mistral-7B-Instruct-v0.3 (Jiang et al. 2023), trained with self-generated data from the ShareGPT dataset (shareAI 2023). Our method is evaluated on LongBench (Bai et al. 2023), RULER (Hsieh et al. 2024), and Needle-in-a-Haystack (Kamradt 2023). Compared to other methods performing KV cache eviction during the pre-filling stage, our approach consistently achieves superior results under the same budget and more closely approaches the theoretical upper bound of using actual decoded tokens.
@@ 2.1 KV Cache Eviction
To address the challenge of reducing the memory footprint of KV cache, a series of methods are proposed from the perspective of reducing the number of elements in KV cache. These strategies can effectively control KV cachen size by evaluating the importance of key-value pairs within the input sequence during pre-filling, retaining only those with high importance scores and discarding less critical ones. This approach is grounded in the observation of the sparsity of attention (Zhang et al. 2023b) and that only a small portion of tokens convey essential information for response generation during the pre-filling stage (Li et al. 2024b). Most of these methods compute the cumulative attention scores of key-value pairs with respect to queries within the local window during the prefilling stage, which are used as the corresponding importance scores for eviction. Methods like H2O (Zhang et al. 2023b) and Scissorhands (Liu et al. 2023) perform dynamic KV cache eviction, while approaches such as SnapKV (Li et al. 2024b) conducts a static eviction. Additionally, some work (Feng et al. 2024; Cai et al. 2024; Yang et al. 2024a) focuses on dynamically allocating budgets to optimize memory usage. Although these methods can achieve effective KV cache eviction during pre-filling stage. As shown in Fig 1,their directly relying on the queries of local window to score the importance of key-value pairs often results in an overemphasis on local information while neglecting global content, thereby limiting performance. Therefore, it is essential to develop a KV cache eviction method that can retain more information during the prefilling stage to support subsequent generation effectively.
@@ 4.2 Training Model With Soft Tokens
Before training, we define the number of soft tokens as $n$ and expand the model’s vocabulary by appending these tokens. During the training process, we exclusively fine-tune the subset of parameters in the model’s embedding layer that correspond to these soft tokens, while keeping the remaining parameters frozen. We construct our training set using 50,000 samples from the ShareGPT dataset, with processing details provided in the appendix. Each sample contains two fields: prompt and response, where the response represents the model’s actual output for the given prompt.
@@ 5 Experiments
Llama-3.1-8B-Instruct
Full KV 31.85 15.59 28.39 29.93 22.98 18.20 34.18 23.79 27.14 72.50 92.14 43.66 8.45 97.59 61.36 51.88 41.23
@@
Mistral-7B-Instruct-v0.3
Full KV 29.20 41.58 53.11 49.27 39.01 28.58 31.95 25.67 23.89 76.00 88.59 47.50 5.50 98.00 51.97 46.28 46.01
@@ 5.2 Main Results
Llama-3.1-8B-Ins
Sequence Length = 8192
@@
Mistral-7B-Ins-v0.3
Sequence Length = 8192
@@
Table 3: Performance comparison of different methods on RULER under different budgets across various LLMs. The number of samples in the test dataset is 100, and all other settings are the same as Table 2.
@@
We evaluate the performance of the methods under different budgets with input lengths of 8192 and 32768, respectively. The experimental results are presented in Table 3. Our method significantly outperforms all baselines, consistently maintaining an advantage of over 3 points and and peaking at nearly 10 points. Additional results under other settings are provided in the appendix.
@@
The experimental results are shown in Figure 4. It is evident that our method significantly exceeds all baselines, indicating its robust performance in retrieval scenarios. Detailed visualization results are available in the appendix.
@@ 6.1 The Effect of Soft Tokens
Figure 4: Performance comparison of different methods on Needle-in-a-Haystack under different budgets across various LLMs. The settings of all methods are the same as the previous experiments. We set the minimum length in the test to 4000, the maximum length to 32000, and the intervals for the document depth percent to 10. The final score is the average of the scores across all test lengths.
@@
(a) Performance of Llama-3.2-1B-Instruct on LongBench across training datasets (Wiki_en = Wikipedia_en, Git = GitHub repos), using consistent setup and a budget of 512.
@@
We further evaluate the effectiveness of Judge Q. Following the prompt template adjustment settings described in Section 3, we test the proportion of performance degradation when the question is placed at the beginning of the prompt. Under the same settings, our method exhibits less than 7% performance loss, which is lower compared to the approximately 10% loss observed in baselines.
@@
Table 5: Perfomance of our method and SnapKV in text continuation tasks, where the KV cache budget is set to 25% of the average output length. Specifically, the budget is set to 1024 for MATH-500 and 3072 for AIME24.
@@ 7 Conclusion
Appendix A Appendix
A.1 Construction of Training Data
As introduced in the main text, we extract 50,000 data entries from ShareGPT for the construction of the training dataset. Each constructed sample contains two fields: prompt and response. To ensure comprehensive coverage of the domain, we extracted 45,000 samples from the common subset and 5,000 from the computer subset. Subsequently, we concatenate the multi-turn dialogues of each sample in sequence to form a complete context of considerable length. This complete dialogue history is truncated by discarding the last 10% content, retaining the initial 90% of the dialogue content as the prompt part in the training data. Then, we continue to generate from this truncated dialogue history using the model to be trained, with the newly generated dialogue content serving as the response. To enable the model to learn information from different positions within the output, we randomly truncate the response. The final data information is as follows, with Llama-3.1-8B-Instruct serving as the model to be trained.
Parameter Value
Training set sample number 50000
Validation set sample number 5000
Avg prompt length(tokens) 959.22
Avg response length(tokens) 85.03
Table 6: Data information.
A.2 Training setups
We utilize a single A100 GPU(40GB) for training and inference. The environment we used includes python=3.10, torch=2.6.0, transformers=4.53.0, and CUDA 12.4. Below are some of the hyperparameter settings used during the training process:
Hyperparameter Value
optimizer AdamW
epoch 3
learning rate 5e-5
batch size 32
Table 7: Hyperparameters for training.
A.3 Prompt adjustion of Single-QA tasks
In Section 3, we modify the prompt template for Single-QA tasks to evaluate the performance of the current pruning method when the question is placed at the beginning. Below are the specific prompt templates before and after the adjustment, the key parts are bolded:
Prompt Adjustment Raw Prompt Template NarrativeQA You are given a story, which can be either a novel or a movie script, and a question. Answer the question asconcisely as you can, using a single phrase if possible. Do not provide any explanation. Story: {context} Now, answer the question based on the story asconcisely as you can, using a single phrase if possible. Do not provide any explanation. Question: {input} Answer: Qasper You are given a scientific article and a question. Answer the question as concisely as you can, using a single phrase or sentence if possible. If the question cannot be answered based on the information in the article, write ”unanswerable”. If the question is a yes/no question, answer ”yes”, ”no”, or ”unanswerable”. Do not provide any explanation. Article: {context} Answer the question based on the above article as concisely as you can, using a single phrase or sentence if possible. If the question cannot be answered based on the information in the article, write ”unanswerable”. If the question is a yes/no question, answer ”yes”, ”no”, or ”unanswerable”. Do not provide any explanation. Question: {input} Answer:
MultiFieldQA-en Read the following text and answer briefly. {context} Now, answer the following question based on the above text, only give me the answer and do not output any other words. Question: {input} Answer:
Reverse Prompt Template NarrativeQA You are given a question, and a story, which can be either a novel or a movie script. Answer the question asconcisely as you can, using a single phrase if possible. Do not provide any explanation. Question: {input}. Story: {context} Now, answer the question based on the story asconcisely as you can, using a single phrase if possible. Do not provide any explanation. Answer: Qasper You are given a question and a scientific article. Answer the question as concisely as you can, using a single phrase or sentence if possible. If the question cannot be answered based on the information in the article, write ”unanswerable”. If the question is a yes/no question, answer ”yes”, ”no”, or ”unanswerable”. Do not provide any explanation. Question: {input} Article: {context} Answer the above question based on the above article as concisely as you can, using a single phrase or sentence if possible. If the question cannot be answered based on the information in the article, write ”unanswerable”. If the question is a yes/no question, answer ”yes”, ”no”, or ”unanswerable”. Do not provide any explanation. Answer: MultiFieldQA-en Read the following question and text, then answer briefly. Question: {input} {context} Now, answer the above question based on the above text, only give me the answer and do not output any other words. Answer:
A.4 Results on RULER
In Section 5.2, we provide the performance results for sequence lengths of 8k and 32k. Additionally, we test the effect of different methods under input length is 16k. The experimental results are shown in Table 8. When the sequence length is 16k, our method consistently outperforms the baselines by at least 5 points across multiple budgets, and in some cases by more than 10 points. All the experimental results strongly demonstrate the effectiveness and generalization of Judge Q.
Budget Method
Full KV StrLLM H2O SnapKV PyraKV Ours
Llama-3.1-8B-Instruct
256 84.40 15.19 24.47 53.28 54.19 59.35
512 84.40 17.16 37.10 59.07 59.45 65.82
1024 84.40 18.64 48.03 64.66 63.26 70.70
Mistral-7B-Ins-v0.3
256 77.92 9.52 12.50 33.33 31.99 43.28
512 77.92 10.94 14.37 43.30 41.98 51.64
1024 77.92 12.90 20.68 50.41 48.36 61.67
Table 8: Performance comparison of different methods on RULER under different budgets across various LLMs when test data sequnece length is 16384. The number of samples in the test dataset is 100, and all other settings are the same as Table 3.
A.5 Results on Needle-in-a-Haystack
In the main text, due to space limitations, we only provide the scores on Needle-in-a-Haystack. Below is a visualization of the performance of our method compared to the best-performing baseline method under different budgets and models. The visualization results show the advantages of our method directly, which further demonstrates the excellent performance of Judge Q in retrieval-related tasks.
Figure 6: Performance of Llama-3.1-8B-Instruct using Judge Q and SnapKV at a budget of 64.
Figure 7: Performance of Llama-3.1-8B-Instruct using Judge Q and SnapKV at a budget of 96.
Figure 8: Performance of Llama-3.1-8B-Instruct using Judge Q and PyramidKV at a budget of 128.
Figure 9: Performance of Mistral-7B-Instruct-v0.3 using Judge Q and SnapKV at a budget of 64.
Figure 10: Performance of Mistral-7B-Instruct-v0.3 using Judge Q and PyramidKV at a budget of 96.
Figure 11: Performance of Mistral-7B-Instruct-v0.3 using Judge Q and SnapKV at a budget of 128.

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable performance across a wide range of domains, including language modeling (Hendrycks et al. 2020; Suzgun et al. 2022), text comprehension (Rajpurkar et al. 2016; Joshi et al. 2017), and code understanding (Chen et al. 2021; Jain et al. 2024). Currently, many scenarios (Liu et al. 2025a, b; Chen et al. 2025; Lu et al. 2025a, b) involve input sequences with over ten thousand tokens, such as books, code repositories (kočiský2017narrativeqareadingcomprehensionchallenge; Zhang et al. 2023a), and dialogue systems with long-term histories (Li et al. 2024a). LLMs employ KV cache to store historical information during the pre-filling stage, and reuse them during decoding. The size of KV cache increases linearly with input sequence length. In long sequence scenarios, KV cache often consumes substantial memory space, posing a significant bottleneck for inference, particularly in resource-limited environments.
@@
This straightforward approach of directly selecting the last window is simple to implement, however, it has limitations. Firstly, it tends to overly focus on local information within the last window while neglecting global content, which can result in potential information loss. Additionally, if the last window does not contain question, the performance of these methods often significantly degrades. In fact, the primary objective of KV cache eviction is to reduce memory usage while maintaining decoding quality. Therefore, identifying the key-value pairs that are crucial for decoding the subsequent tokens is logical and ideal.
In this paper, we discover that directly using the original decoded tokens to select critical key-value pairs preserves high decoding quality, representing the theoretical upper bound of KV cache eviction methods. However, during the prefilling stage, the actual decoded tokens are not available. Inspired by this phenomenon, we propose a training method called Judge Q. This approach involves introducing a list of soft tokens into the model’s vocabulary and leveraging them to perform KV cache eviction during the pre-filling stage. Specifically, during the training phase, we concatenate these soft tokens directly to the end of original input sequence. The soft tokens are trained to align the attention map to input sequence with that of the actual decoded tokens, enabling the soft tokens to capture global information effectively. This process incurs minimal training costs, as it involves training only the parameters associated with the soft tokens in the model’s embedding layer, while keeping the rest of the model’s weights frozen. During inference, we similarly append these soft tokens to the end of the original input sequence in the pre-filling stage. The queries corresponding to soft tokens are used to calculate the importance scores of the key-value pairs in KV cache across the entire input sequence. Based on these scores, we retain the top-k key-value pairs with the highest scores and discard the rest. After the pre-filling stage, the soft tokens are also removed. The pruned KV cache is then used for the subsequent decoding process. We conduct experiments using two models, Llama-3.1-8B-Instruct (Dubey et al. 2024) and Mistral-7B-Instruct-v0.3 (Jiang et al. 2023), trained with self-generated data from the ShareGPT dataset (shareAI 2023). Our method is evaluated on LongBench (Bai et al. 2023), RULER (Hsieh et al. 2024), and Needle-in-a-Haystack (Kamradt 2023). Compared to other methods performing KV cache eviction during the pre-filling stage, our approach consistently achieves superior results under the same budget and more closely approaches the theoretical upper bound of using actual decoded tokens.
@@ 2.1 KV Cache Eviction
To address the challenge of reducing the memory footprint of KV cache, a series of methods are proposed from the perspective of reducing the number of elements in KV cache. These strategies can effectively control KV cache size by evaluating the importance of key-value pairs within the input sequence during pre-filling, retaining only those with high importance scores and discarding less critical ones. This approach is grounded in the observation of the sparsity of attention (Zhang et al. 2023b) and that only a small portion of tokens convey essential information for response generation during the pre-filling stage (Li et al. 2024b). Most of these methods compute the cumulative attention scores of key-value pairs with respect to queries within the local window during the prefilling stage, which are used as the corresponding importance scores for eviction. Methods like H2O (Zhang et al. 2023b) and Scissorhands (Liu et al. 2023) perform dynamic KV cache eviction, while approaches  (Li et al. 2024b; Chen et al. 2024; He et al. 2025) conduct a static eviction. Additionally, some work (Feng et al. 2024; Cai et al. 2024; Yang et al. 2024a) focuses on dynamically allocating budgets to optimize memory usage. Although these methods achieve effective KV cache eviction during pre-filling stage. As shown in Fig 1,their directly relying on the queries of local window to score the importance of key-value pairs often results in an overemphasis on local information while neglecting global content, thereby limiting performance. Therefore, it is essential to develop a KV cache eviction method that can retain more information during the prefilling stage to support subsequent generation effectively.
@@ 4.2 Training Model With Soft Tokens
Before training, we define the number of soft tokens as $n$ and expand the model’s vocabulary by appending these tokens. During the training process, we exclusively fine-tune the subset of parameters in the model’s embedding layer that correspond to these soft tokens, while keeping the remaining parameters frozen. We construct our training set using 50,000 samples from the ShareGPT dataset, with processing details provided in the repository. Each sample contains two fields: prompt and response, where the response represents the model’s actual output for the given prompt.
@@ 5 Experiments
Llama-3.1-8B-Instruct Full KV 31.85 15.59 28.39 29.93 22.98 18.20 34.18 23.79 27.14 72.50 92.14 43.66 8.45 97.59 61.36 51.88 41.23
@@
Mistral-7B-Instruct-v0.3 Full KV 29.20 41.58 53.11 49.27 39.01 28.58 31.95 25.67 23.89 76.00 88.59 47.50 5.50 98.00 51.97 46.28 46.01
@@ 5.2 Main Results
Llama-3.1-8B-Ins Sequence Length = 8192
@@
Mistral-7B-Ins-v0.3 Sequence Length = 8192
@@
Table 3: Performance on RULER under different budgets across various LLMs. The number of samples in the test dataset is 100, and all other settings are the same as Table 2.
@@
We evaluate the performance of the methods under different budgets with input lengths of 8192 and 32768, respectively. The experimental results are presented in Table 3. Our method significantly outperforms all baselines, consistently maintaining an advantage of over 3 points and and peaking at nearly 10 points. Additional results under other settings are provided in the repository.
@@
The experimental results are shown in Figure 4. It is evident that our method significantly exceeds all baselines, indicating its robust performance in retrieval scenarios. Detailed visualization results are available in the repository.
@@ 6.1 The Effect of Soft Tokens
Figure 4: Performance on Needle-in-a-Haystack. The settings are the same as previous experiments. We set the minimum length to 4000, the maximum length to 32000, and the intervals for the document depth percent to 10.
@@
(a) Performance of Llama-3.2-1B-Instruct on LongBench across training datasets (SG = ShareGPT, Wiki = Wikipedia_en, Git = GitHub repos), using consistent setup and a budget of 512.
@@
We further evaluate the effectiveness of Judge Q. Following the prompt template adjustment settings described in Section 3, we test the proportion of performance degradation when the question is placed at the start of the prompt. Under the same settings, our method exhibits less than 7% performance loss, which is lower compared to the approximately 10% loss observed in baselines.
@@
Table 5: Performance on text continuation tasks, under a budget of 25% of average output length. Specifically, the budget is set to 1024 for MATH-500 and 3072 for AIME24.
@@ 7 Conclusion
Acknowledgments
We gratefully acknowledge the support of the National Natural Science Foundation of China (NSFC) via grant 62236004, 62206078 and 62476073.